For the following pairs of lines, determine whether the lines are parallel, perpendicular or neither

3x-9y=6 and 1x-7y=7
Answer #1

Solution. 32-gu-6. and 12- 74=7. It may write as. 94 32-6 y = 13x - 2 - compairing with y = mpete, m₂ = 1 and 12-74=7 Lorite